OpenStudy (anonymous):

Identify the slope and y-intercept of the graph of the equation. Then graph the equation. y = -(5/4) x + 1 A. slope: -(5/4) y-intercept: –1

OpenStudy (anonymous):

A. -(5/4) y- intercept: -1 http://i1188.photobucket.com/albums/z417/esmy_071/mc009-3.jpg

OpenStudy (anonymous):

B. -(4/5) Y-intercept: 1 http://i1188.photobucket.com/albums/z417/esmy_071/mc019-1.jpg

OpenStudy (jackellyn):

y = mx +b m = slope b= y intercept rise slope =---- run

OpenStudy (anonymous):

you got the slope right

OpenStudy (jackellyn):

y intercept, on the graph, will be where the line crosses the y axis.

OpenStudy (jackellyn):

@satellite73 i think she is putting the choices, that isn't her answer.

OpenStudy (anonymous):

i am confused by your question are you supposed to graph \(y=-\frac{5}{4}x+1\) or are you supposed to find the equation of the line graphed above

OpenStudy (anonymous):

oh then let me be quiet

OpenStudy (anonymous):

Identify the slope and y-intercept of the graph of the equation. Then graph the equation. y = -(5/4) x + 1

OpenStudy (anonymous):

I think its a or b :/

OpenStudy (anonymous):

But ill put the whole thing to see if I'm wrong

OpenStudy (anonymous):

ooh i see you have the slope correct, it is \(-\frac{5}{4}\) and the y - intercept is 1

OpenStudy (anonymous):

which means it crosses the y axis at \((0,1)\) so it is your second graph above

ganeshie8 (ganeshie8):

no. both a & b are not matching

OpenStudy (anonymous):

nope scratch taht!!

OpenStudy (anonymous):

it is not the second graph, because although the y intercept is one, the slope is not \(-\frac{5}{4}\)

OpenStudy (anonymous):

thanks to @ganeshie8 for catching that one

OpenStudy (anonymous):

C. slope: -(5/4) y-intercept: 1 http://i1188.photobucket.com/albums/z417/esmy_071/mc009-7.jpg

OpenStudy (anonymous):

Its C

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

you are right its C. good work :)

OpenStudy (anonymous):

C matches the slope is write in C and it lines up good. Thank you :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

yes it is C